OpenStudy (anonymous):

Find the exact value of the volume of a can of asparagus with a diameter 6 centimeters and height of 10 centimeters. Include correct units with your solution. Use "pi" in place of the pi symbol.

OpenStudy (saifoo.khan):

pi r^2 x h

OpenStudy (saifoo.khan):

this is the formula to find the volume of a cylinder.

OpenStudy (saifoo.khan):

insert the value & try finding the answer by yourself.

OpenStudy (anonymous):

i dont know anything abt this work is why im on here so how can i find myself

OpenStudy (anonymous):

n pls go view my first question u gave me formula is that correct for my answer

OpenStudy (anonymous):

V = pi r^2 h V = (3.14) (3)^2(10) V = 282.2 cm^3